OpenStudy (anonymous):

help please (: F equals 27 point 2 times m all over t F is the force m is the mass t is the time in seconds Solve F equals 27 point 2 times m all over t for m. Please show your work. (6 points) Use your new formula to find the mass of the automobile in kg for each value of force and time given in the table below. Round your answer to the nearest whole number. (4 points) Force (Newtons) Mass (kg) Time (sec) 11000 8 9520 10 12000 7 11500 9

OpenStudy (jamesj):

So you're saying \[ F = \frac{27.2 m}{t} \] What then are you asking in the first part of the question?

OpenStudy (jamesj):

I think you are missing part of the question

OpenStudy (anonymous):

@JamesJ yeah that's right and i have to find out the mass using that formula

OpenStudy (jamesj):

Ah you need to solve for m. Fine. First, multiply both sides by t and you get \[ Ft = 27.2 m \] Can you solve for m now?

OpenStudy (anonymous):

@JamesJ umm i think so?

OpenStudy (jamesj):

Take Ft = 27.2m and divide both sides by 27.2 and you get ... what?

OpenStudy (anonymous):

@JamesJ ummm ft=m?

OpenStudy (jamesj):

That can't be right. You didn't divide the left hand side by 27.2 as well. So, we have F = 27.2m/t Ft = 27.2m multiplying both sides by t Ft/27.2 = m dividing both sides by 27.2 Hence \[ m = \frac{Ft}{27.2} \]

OpenStudy (jamesj):

Now apply that formula for the second part of the question. You are given values of F and t. Use this formula to find the corresponding values of m.

OpenStudy (anonymous):

@JamesJ ohh okay i think i get it now